# ShapeMoiré: Channel-Wise Shape-Guided Network for Image Demoiréing

### v1

@@
Abstract
Photographing optoelectronic displays often introduces unwanted moiré patterns due to analog signal interference between the pixel grids of the display and the camera sensor arrays. This work identifies two problems that are largely ignored by existing image demoiréing approaches: 1) moiré patterns vary across different channels (RGB); 2) repetitive patterns are constantly observed. However, employing conventional convolutional (CNN) layers cannot address these problems. Instead, this paper presents the use of our recently proposed Shape concept. It was originally employed to model consistent features from fragmented regions, particularly when identical or similar objects coexist in an RGB-D image. Interestingly, we find that the Shape information effectively captures the moiré patterns in artifact images. Motivated by this discovery, we propose a ShapeMoiré method to aid in image demoiréing. Beyond modeling shape features at the patch-level, we further extend this to the global image-level and design a novel Shape-Architecture. Consequently, our proposed method, equipped with both ShapeConv and Shape-Architecture, can be seamlessly integrated into existing approaches without introducing additional parameters or computation overhead during inference. We conduct extensive experiments on four widely used datasets, and the results demonstrate that our ShapeMoiré achieves state-of-the-art performance, particularly in terms of the PSNR metric. We then apply our method across four popular architectures to showcase its generalization capabilities. Moreover, our ShapeMoiré is robust and viable under real-world demoiréing scenarios involving smartphone photographs. We open-sourced an implementation of ShapeMoiré in PyTorch at https://github.com/ShapeMoire.
Index Terms:
@@ Abstract
I Introduction
@@
Moiré artifacts appear as wavy or rippled distortions in digital photographs. This unwanted pattern is caused by the interference between the color filter array of cameras and the sub-pixel arrangement of displays. Fig. 1 illustrates a digital photo with such moiré patterns ($I_{moir\acute{e}}$), which make the original image ($I$) appear opaque and distorted when displayed on a screen. Notably, the irregular shape and wide frequency span of moiré patterns distinguish demoiréing from other image restoration tasks such as super-resolution [1, 2] and reducing JPEG compression artifacts [3].
@@
Figure 1: Visual demonstration of the moiré pattern in the given image and its separate manifestations in the individual RGB channels. Left column: the original image $I$, where the color in each channel is uniformly and smoothly distributed; Middle column: it can be observed that the three channels of $I_{moir\acute{e}}$ display distinct moiré patterns; Right column: the shapes of moiré patterns in the shape space ($R^{Shape}_{moir\acute{e}}$, $G^{Shape}_{moir\acute{e}}$ and $B^{Shape}_{moir\acute{e}}$) remain consistent with the patterns in the different channels of the original moiré image ($R_{moir\acute{e}}$, $G_{moir\acute{e}}$ and $B_{moir\acute{e}}$), but with enhanced structural integrity.
@@
Moiré pattern removal has attracted extensive research interests over the past few years [4, 5, 6, 7, 8]. Initially, efforts were devoted to collecting synthetic data [9] or genuine images with low resolution [10]. Subsequent datasets expanded the image resolution to 1080P [6] or even ultra-high-definition (4K) [7], posing greater challenges in removing moirés without compromising the pristine features of the images. To address this, prior methods have mostly relied on rich side information, such as complementary visual attributes [5] and aliasing frequencies [11, 12]. Despite the advancements in existing methods, they often treat the input image from a holistic perspective, especially without discriminating between different channels (i.e., RGB channels). However, as observed in Fig. 1, unlike the uniform and smooth color distribution in each channel ($R$, $G$ and $B$) of the original image $I$, these three input channels ($R_{moir\acute{e}}$, $G_{moir\acute{e}}$ and $B_{moir\acute{e}}$) typically exhibit distinct moiré patterns. As a result, processing them equally with convolutional layers confuses the model’s focus on demoiréing, resulting in sub-optimal outcomes. Furthermore, existing literature often overlooks the repetitive patterns that are consistently observed.
@@
Our idea in this work is inspired by a recent prevailing concept—Shape—in RGB-D segmentation [13]. Unlike traditional CNNs, ShapeConv [13] can distinguish between different channels and is adept at modeling repetitive shapes, such as two identical chairs in a depth image from RGB-D segmentation. In the context of moiré patterns in an image, as demonstrated in Fig. 1, the moiré image ($I_{moir\acute{e}}$) exhibits frequent similar or repetitive patterns, unlike the uniform color distribution in the original image ($I$). This type of information is precisely what shape excels at expressing. Additionally, the features of moiré patterns in the shape space ($R^{Shape}_{moir\acute{e}}$, $G^{Shape}_{moir\acute{e}}$ and $B^{Shape}_{moir\acute{e}}$) remain consistent across different channels of the original moiré image ($R_{moir\acute{e}}$, $G_{moir\acute{e}}$ and $B_{moir\acute{e}}$), but with enhanced structural integrity. This result motivates us to approach the image demoiré problem from the shape perspective.
@@
Therefore, we propose a novel method named ShapeMoiré to extend our ShapeConv for image demoiring. Our ShapeMoiré operates on both patch [^1] The operational unit of input features for the convolutional layer, whose spatial size matches that of the convolution kernel.-level and image-level shapes. This results in two key components of our proposed ShapeMoiré: the Shape-aware convolutional layer (ShapeConv) and Shape-aware network architecture (Shape-Architecture). Specifically, at the patch-level, we first decompose a patch into two separate components, i.e., a base-component and a shape-component. We then employ two operations, namely, base-product and shape-product, to process these components respectively with two learnable weights: the base-kernel and the shape-kernel. The output from these operations are then combined in an addition manner to form a shape-aware patch, which is subsequently convolved with a standard convolutional kernel. Unlike the original patch, the shape-aware patch is capable of adaptively learning the shape characteristic with the shape-kernel, while the base-kernel serves to balance the contributions of the shape and the base components in the final feature representation. For the extended image-level, we incorporate an additional training stream to enhance the model’s focus on shape information during training. This is achieved by adding a loss function for the shape branch, thereby updating the parameters with an emphasis on shape information. In a nutshell, the former replaces the conventional convolutional layer with a delicately designed ShapeConv that involves the modeling of shape features, and the latter employs the shape information of raw pixels to further enhance the baseline architecture. It is worth noting that our ShapeMoiré can be seamlessly integrated into most baseline architectures that are equipped with convolutional layers as building blocks. In addition, benefiting from its original virtue, our ShapeMoiré results in no incremental parameters and very negligible computational cost.
@@
To evaluate the effectiveness of our proposed method, we conducted extensive experiments on four well-established datasets: LCDMoiré [9], TIP2018 [10], FHDMi [6], and UHDM [7]. Our experimental results demonstrate that ShapeMoiré outperforms existing state-of-the-arts by a significant performance margin. To assess its generalization capabilities, we applied ShapeMoiré to various widely adopted demoiréing architectures [10, 11, 7] and observed improved results compared to baselines. Notably, qualitative results highlight that ShapeMoiré remarkably enhances moiré pattern removal, even when applied to real-world images captured using a smartphone. This underscores both the effectiveness on public datasets and the practical applicability of our ShapeMoiré in real-world scenarios.
@@
The work on ShapeConv has previously been published at ICCV [13]. This extended manuscript expands upon the original version in several key aspects:
@@
  • We broaden the application of the notion of Shape beyond its original use in RGB-D segmentation. For the first time, this concept has been effectively applied to the task of removing moiré pattern in 2D images.
@@
  • Beyond employing patch-level ShapeConv, we introduce an image-level Shape-Architecture for image demoiréing. Through ablation studies, we have demonstrated that the Shape information at different levels can enhance performance on this challenging task.
@@
  • The proposed ShapeMoiré method, which incorporates both ShapeConv and Shape-Architecture, represents a versatile approach. We conduct experiments using various existing methods as baselines, all of which effectively enhance baseline performance. Notably, both ShapeConv and Shape-Architecture introduce ZERO additional parameters and negligible computational overhead during inference.
II Related Work
II-A Image Demoiréing
@@
Removing moiré patterns constitutes a form of image denoising aimed at mitigating the destructive artifacts arising from frequency interference during image capture. Unlike other types of noise, such as flickering [14] induced by photography, there is currently no standard tool for effectively eliminating moiré patterns. Recently, numerous benchmark datasets and approaches have been introduced to address the challenge of image demoiréing.
@@
Liu et al. [15] initiated the task by curating a synthetic dataset that simulates the process of photographing with a camera. Building upon this approach, Yuan et al. [9] developed a large-scale synthetic dataset, namely, LCDMoiré, for further image demoiréing [9, 12, 16]. However, the transition from synthetic training to real-world scenarios presents a significant obstacle due to the inherent gap between simulation and reality. Sun et al. [10] addressed this challenge by introducing the TIP2018 dataset, which utilizes real-world moiré images. Following this approach, Yue et al.[8] compiled a well-aligned dataset of raw moiré image. Nevertheless, these images often have low resolutions, making it difficult for the accompanying methods to adapt to high-resolution images. To tackle this issue, He et al. [6] curated the FHDMi dataset with a resolution of 1080P. However, as the demand for image quality continues to increase, the practicality of 1080p resolution may diminish, especially when considering the ultra-high-definition (4K) images captured by modern mobile cameras. Recognizing these challenges and limitations, Yu et al. [7] collected the first ultra-high-definition demoiréing dataset (UHDM), which contains 5,000 real-world 4K resolution image.
@@ Methods
Unlike previous methods, which treat the entire image as a whole without distinguishing between different channels, Yang et al. [4] also observed differences in moiré patterns across different RGB channels. Specifically, they found that the red (R) and blue (B) channels are more heavily affected by moiré artifacts compared to the green (G) channel. Therefore, they proposed to remove moiré artifacts in the R and B channels using guided filtering based on the texture layer obtained from the G channel. In contrast to the method in [4], our approach does not involve special treatment of a specific channel but instead presents a universal method that discriminates between different channels. Moreover, [4] is a specialized design, while our proposed ShapeMoiré stands out as a versatile approach. It can seamlessly integrate with existing methods, enhancing model performance with ZERO increase of memory and computation.
II-B Image Restoration
@@
Image restoration aims to reconstruct the original, clean and high-quality image from degraded versions. This task encompasses various processes, including image super-resolution (SR), image denoising, reducing JPEG compression artifacts and so on. A plethora of learning-based image restoration models [17, 18, 19, 18], have gained popularity due to their impressive performance. These models typically leverage large-scale paired datasets to learn mappings between low-quality and high-quality images, employing encoder-decoder structures [20] or hierarchical architectures [10, 21]. Building upon pioneering works such ase SRCNN [22] (for SR), DnCNN [23] (for denoising), and ARCNN [3] (for reducing JPEG artifacts), numerous CNN-based models [24, 25, 26] have emerged. These models enhance representational capacity by incorporating larger and deeper neural architectures, such as residual blocks [27], dense blocks [28], and others [24].
II-C Shape Concept and Its Applications
@@
The concept of Shape was initially introduced in the domain of RGB-D segmentation in a previously published paper [13], where it is defined as the disparity between local depth values and the overall distance of the region from the observation point. One key observation from this paper is that, in spaces where identical or similar objects coexist, Shape can extract more consistent features compared to raw pixel values. To leverage this insight, we designed ShapeConv to replace traditional convolution layers, allowing for the processing of data from various modalities while emphasizing Shape attributes. Subsequent studies  [29, 30] extended the application to remote sensing data, showcasing Shape’s superiority in feature extraction across RGB+nDSM (primarily representing height features) data modalities. Specifically, Xiong et al. [29] utilized ShapeConv for semantic segmentation tasks on the their collected GAMUS dataset, which features numerous repetitive buildings. Zhong et al. [30] created an ocean coral dataset characterized by repetitive texture patterns within the corals. By integrating ShapeConv as a fundamental component of their model, they achieved superior performance. Beyond the RGB+X data, [31] employs ShapeConv for 3D object detection and occupancy.
@@
III Method
Figure 2: Overview of ShapeMoiré. The top part depicts the network’s overall architecture. The green and blue parts denote the baseline and ShapeMoiré components, respectively. and The red line flow represents ShapeMoiré during inference. The bottom part shows network layers: (a) Convolution-Layer (consistent in both training and testing), and (b) and (c) illustrate ShapeConv during training and inference, respectively. For both ShapeConv and Shape-Architecture, our method introduces no parameters and negligible computational cost during inference.

### v2

@@
Abstract.
Photographing optoelectronic displays often introduces unwanted moiré patterns due to analog signal interference between the pixel grids of the display and the camera sensor arrays. This work identifies two problems that are largely ignored by existing image demoiréing approaches: 1) moiré patterns vary across different channels (RGB); 2) repetitive patterns are constantly observed. However, employing conventional convolutional (CNN) layers cannot address these problems. Instead, this paper presents the use of our recently proposed Shape concept. It was originally employed to model consistent features from fragmented regions, particularly when identical or similar objects coexist in an RGB-D image. Interestingly, we find that the Shape information effectively captures the moiré patterns in artifact images. Motivated by this discovery, we propose a new method, ShapeMoiré, for image demoiréing. Beyond modeling shape features at the patch-level, we further extend this to the global image-level and design a novel Shape-Architecture. Consequently, our proposed method, equipped with both ShapeConv and Shape-Architecture, can be seamlessly integrated into existing approaches without introducing any additional parameters or computation overhead during inference. We conduct extensive experiments on four widely used datasets, and the results demonstrate that our ShapeMoiré achieves state-of-the-art performance, particularly in terms of the PSNR metric. We then apply our method across four popular architectures to showcase its generalization capabilities. Moreover, to further validate its generality beyond the demoiréing task, we apply ShapeMoiré to the image deblurring task, where it continues to deliver consistent performance gains. Finally, experiments on real-world images captured by smartphones confirm the robustness and practical applicability of ShapeMoiré in challenging demoiréing scenarios. We open-sourced an implementation of ShapeMoiré in PyTorch at https://github.com/ShapeMoire.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Machine learning approaches
1. Introduction
@@
Moiré artifacts manifest as wavy or rippled distortions in digital photographs, arising from the interference between a camera’s color filter array and the sub-pixel structure of electronic displays. These unwanted patterns significantly degrade visual quality, posing challenges in multimedia applications such as mobile photography (Rawat and Kankanhalli, 2015) and cross-device content reproduction (Ng et al., 2017). Fig. 1 illustrates a digital photo with such moiré patterns ($I_{moir\acute{e}}$), which make the original image ($I$) appear opaque and distorted when displayed on a screen. Notably, the irregular shape and broad frequency ranges of moiré patterns distinguish demoiréing from other image restoration tasks such as super-resolution (Zhang et al., 2015; Wei et al., 2023) and reducing JPEG compression artifacts (Dong et al., 2015).
@@
Figure 1. Visual demonstration of the moiré pattern in the given image and its separate manifestations in the individual RGB channels. Left column: the original image $I$, where the color in each channel is uniformly and smoothly distributed; Middle column: it can be observed that the three channels of $I_{moir\acute{e}}$ display distinct moiré patterns; Right column: the shapes of moiré patterns in the shape space ($R^{Shape}_{moir\acute{e}}$, $G^{Shape}_{moir\acute{e}}$ and $B^{Shape}_{moir\acute{e}}$) remain consistent with the patterns in the different channels of the original moiré image ($R_{moir\acute{e}}$, $G_{moir\acute{e}}$ and $B_{moir\acute{e}}$), but with enhanced structural integrity.
@@
Moiré pattern removal has attracted extensive research interests over the past few years (Yang et al., 2017; He et al., 2019, 2020; Yu et al., 2022; Yue et al., 2022). Initially, efforts were devoted to collecting synthetic data (Yuan et al., 2019) or genuine images with low resolution (Sun et al., 2018). Subsequent datasets expanded the image resolution to 1080P (He et al., 2020) or even ultra-high-definition (4K) (Yu et al., 2022), posing greater challenges in removing moirés without compromising the pristine features of the images. To address this, prior methods have mostly relied on rich side information, such as complementary visual attributes (He et al., 2019) and aliasing frequencies (Liu et al., 2020; Zheng et al., 2020). Despite the advancements in existing methods, they often treat the input image from a holistic perspective, especially without discriminating between different channels (i.e., RGB channels). However, as observed in Fig. 1, unlike the uniform and smooth color distribution in each channel ($R$, $G$ and $B$) of the original image $I$, these three input channels ($R_{moir\acute{e}}$, $G_{moir\acute{e}}$ and $B_{moir\acute{e}}$) typically exhibit distinct moiré patterns. As a result, processing them equally with convolutional layers confuses the model’s focus on demoiréing, resulting in sub-optimal outcomes. Furthermore, existing literature often overlooks the repetitive patterns that are consistently observed.
@@
Our idea in this work is inspired by a recent prevailing concept—Shape—in RGB-D segmentation (Cao et al., 2021). Unlike traditional CNNs, ShapeConv (Cao et al., 2021) can distinguish between different channels and is adept at modeling repetitive shapes, such as two identical chairs in a depth image from RGB-D segmentation. In the context of moiré patterns in an image, as demonstrated in Fig. 1, the moiré image ($I_{moir\acute{e}}$) exhibits frequent similar or repetitive patterns, unlike the uniform color distribution in the original image ($I$). This type of information is precisely what shape excels at expressing. Additionally, the features of moiré patterns in the shape space ($R^{Shape}_{moir\acute{e}}$, $G^{Shape}_{moir\acute{e}}$ and $B^{Shape}_{moir\acute{e}}$) remain consistent across different channels of the original moiré image ($R_{moir\acute{e}}$, $G_{moir\acute{e}}$ and $B_{moir\acute{e}}$), but exhibit enhanced structural integrity. This observation motivates us to approach the image demoiré problem from the shape perspective.
@@
Therefore, we propose a novel method named ShapeMoiré to extend our ShapeConv for image demoiring. Our ShapeMoiré operates on both the patch [^1] The operational unit of input features for the convolutional layer, whose spatial size matches that of the convolution kernel.-level and image-level shapes. This results in two key components of our proposed ShapeMoiré: the Shape-aware convolutional layer (ShapeConv) and the Shape-aware network architecture (Shape-Architecture). Specifically, at the patch-level, we first decompose a patch into two separate components, i.e., a base-component and a shape-component. We then employ two operations, namely, base-product and shape-product, to process these components separately, with two learnable weights: the base-kernel and the shape-kernel. The outputs from these operations are then combined in an additive manner to form a shape-aware patch, which is subsequently convolved with a standard convolutional kernel. Unlike the original patch, the shape-aware patch is capable of adaptively learning the shape characteristics with the shape-kernel, while the base-kernel serves to balance the contributions of the shape and the base components in the final feature representation. At the extended image-level, we incorporate an additional training stream to enhance the model’s focus on shape information during training. This is achieved by adding a loss function for the shape branch, thereby updating the parameters with an emphasis on shape information. In a nutshell, the former replaces the conventional convolutional layer with a delicately designed ShapeConv that involves the modeling of shape features, and the latter employs the shape information of raw pixels to further enhance the baseline architecture. It is worth noting that our ShapeMoiré can be seamlessly integrated into most baseline architectures that are equipped with convolutional layers as building blocks. In addition, benefiting from its original virtue, our ShapeMoiré results in no incremental parameters and only negligible computational cost.
@@
To evaluate the effectiveness of our proposed method, we conducted extensive experiments on four well-established datasets: LCDMoiré (Yuan et al., 2019), TIP2018 (Sun et al., 2018), FHDMi (He et al., 2020), and UHDM (Yu et al., 2022). Our experimental results demonstrate that ShapeMoiré consistently outperforms existing state-of-the-art methods by a significant margin. To assess its generalization capabilities, we applied ShapeMoiré to various widely adopted demoiréing architectures (Sun et al., 2018; Liu et al., 2020; Yu et al., 2022) and observed improved results compared to baselines. Furthermore, to validate the method’s broader applicability beyond demoiréing, we extended our experiments to an additional image restoration task—image deblurring. Results on the DPDD dataset (Abuolaim and Brown, 2020) confirm that ShapeMoiré remains effective in this new context, indicating good cross-task generalization. Notably, qualitative results highlight that ShapeMoiré remarkably enhances moiré pattern removal, even when applied to real-world images captured using a smartphone. This underscores both the effectiveness on public datasets and the practical applicability of our ShapeMoiré in real-world scenarios.
@@
The work on ShapeConv has previously been published at ICCV (Cao et al., 2021). This extended manuscript expands upon the original version in several key aspects:
@@
  • We broaden the application of the notion of Shape beyond its original use in RGB-D segmentation. To the best of our knowledge, this is the first time the concept has been effectively applied to both the moiré pattern removal and image deblurring tasks in 2D image restoration.
@@
  • Beyond employing patch-level ShapeConv, we introduce an image-level Shape-Architecture for image demoiréing. Through comprehensive ablation studies and cross-task experiments, we demonstrate that integrating Shape information at multiple levels consistently enhances performance on diverse image restoration tasks.
@@
  • The proposed ShapeMoiré method, which incorporates both ShapeConv and Shape-Architecture, represents a versatile approach. We conduct experiments using various existing methods as baselines, all of which effectively enhance baseline performance. Notably, both ShapeConv and Shape-Architecture introduce ZERO additional parameters and incur only negligible computational overhead during inference.
2. Related Work
2.1. Image Demoiréing
@@
Removing moiré patterns constitutes a form of image denoising aimed at mitigating the destructive artifacts arising from frequency interference during image capture. Unlike other types of noise, such as flickering [^2] https://digitalanarchy.com/Flicker/main.html induced by photography, there is currently no standardized tool for effectively eliminating moiré patterns. Recently, numerous benchmark datasets and approaches have been introduced to address the challenge of image demoiréing.
@@
Liu et al. (Liu et al., 2018) initiated the task by curating a synthetic dataset that simulates the process of photographing with a camera. Building upon this approach, Yuan et al. (Yuan et al., 2019) developed a large-scale synthetic dataset, namely, LCDMoiré, for further image demoiréing (Yuan et al., 2019; Zheng et al., 2020; Cheng et al., 2019). However, the transition from synthetic training to real-world scenarios presents a significant obstacle due to the inherent gap between simulation and reality. Sun et al. (Sun et al., 2018) addressed this challenge by introducing the TIP2018 dataset, which utilizes real-world moiré images. Following this approach, Yue et al.(Yue et al., 2022) compiled a well-aligned dataset of raw moiré image. Nevertheless, these images often have low resolutions, making it difficult for the accompanying methods to adapt to high-resolution images. To tackle this issue, He et al. (He et al., 2020) curated the FHDMi dataset with a resolution of 1080P. However, as the demand for image quality continues to increase, the practicality of 1080p resolution may diminish, especially when considering the ultra-high-definition (4K) images captured by modern mobile cameras. Recognizing these challenges and limitations, Yu et al. (Yu et al., 2022) collected the first ultra-high-definition demoiréing dataset (UHDM), which contains 5,000 real-world 4K resolution image.
@@ Methods
Unlike previous methods, which treat the entire image as a whole without distinguishing between different channels, Yang et al. (Yang et al., 2017) also observed differences in moiré patterns across different RGB channels. Specifically, they found that the red (R) and blue (B) channels are more heavily affected by moiré artifacts compared to the green (G) channel. Therefore, they proposed to remove moiré artifacts in the R and B channels using guided filtering based on the texture layer obtained from the G channel. In contrast to the method in (Yang et al., 2017), our approach does not involve special treatment of a specific channel but instead presents a universal method that discriminates between different channels. Moreover, (Yang et al., 2017) is a specialized design, while our proposed ShapeMoiré stands out as a versatile approach. It can seamlessly integrate with existing methods, enhancing model performance with ZERO increase of memory and computation.
2.2. Image Restoration
@@
Image restoration aims to reconstruct the original, clean and high-quality image from degraded versions, serving as a core task in computer vision and multimedia computing. This task encompasses various processes, including image super-resolution (SR) (Jiang et al., 2016), image denoising (Chen and Huang, 2015; Yan et al., 2020; Xu et al., 2023a), reducing JPEG compression artifacts and so on. A plethora of learning-based image restoration models (Liang et al., 2021; Zamir et al., 2021, 2021), have gained popularity due to their impressive performance. These models typically leverage large-scale paired datasets to learn mappings between low-quality and high-quality images, employing encoder-decoder structures (Ronneberger et al., 2015) or hierarchical architectures (Sun et al., 2018; Zhang et al., 2019). Building upon pioneering works such ase SRCNN (Dong et al., 2014) (for SR), DnCNN (Zhang et al., 2017) (for denoising), and ARCNN (Dong et al., 2015) (for reducing JPEG artifacts), numerous CNN-based models (Lai et al., 2017; Zhang et al., 2018b; Wang et al., 2019; Guo et al., 2023) have emerged. These models enhance representational capacity by incorporating larger and deeper neural architectures, such as residual blocks (He et al., 2016), dense blocks (Huang et al., 2017), and others (Lai et al., 2017). In recent years, transformer-based and attention-driven frameworks have further advanced the field. Restormer (Zamir et al., 2022a) proposes a multi-Dconv head transposed attention mechanism tailored for image restoration, achieving state-of-the-art results on multiple benchmarks. Uformer (Wang et al., 2022) unifies transformer and U-Net architectures to improve global context modeling while maintaining efficiency. Guo et al. (Guo et al., 2024) introduces a simple but effective baseline, named MambaIR, which introduces both local enhancement and channel attention, yet surprisingly achieves competitive performance on several restoration tasks. Lin et al. (Lin et al., 2024) maps the degraded images into textual representations for removing the degradations, and then convert the restored textual representations into a guidance image for assisting image restoration.
However, while these methods target general image restoration tasks—such as image super-resolution, image denoising, and JPEG compression artifact reduction—the denoising types they address are mostly spatial-domain noises like Gaussian color image denoising (Zamir et al., 2022a; Guo et al., 2024), real image denoising (Guo et al., 2024), deraining (Wang et al., 2022), and deblurring (Lin et al., 2024). Although moiré patterns are also a type of noise, they are frequency-domain artifacts with unique structural characteristics and thus have not been effectively tackled by these existing methods.
2.3. Shape Concept and Its Applications
@@
The concept of Shape was initially introduced in the domain of RGB-D segmentation in a previously published paper (Cao et al., 2021), where it is defined as the disparity between local depth values and the overall distance of the region from the observation point. One key observation from this paper is that, in spaces where identical or similar objects coexist, Shape can extract more consistent features compared to raw pixel values. To leverage this insight, we designed ShapeConv to replace traditional convolution layers, allowing for the processing of data from various modalities while emphasizing Shape attributes. Subsequent studies  (Xiong et al., 2023; Zhong et al., 2023) extended the application to remote sensing data, showcasing Shape’s superiority in feature extraction across RGB+nDSM (primarily representing height features) data modalities. Specifically, Xiong et al. (Xiong et al., 2023) utilized ShapeConv for semantic segmentation tasks on the their collected GAMUS dataset, which features numerous repetitive buildings. Zhong et al. (Zhong et al., 2023) created an ocean coral dataset characterized by repetitive texture patterns within the corals. By integrating ShapeConv as a fundamental component of their model, they achieved superior performance. Beyond the RGB+X data, (Zhou et al., 2024) employs ShapeConv for 3D object detection and occupancy.
@@
3. Method
Figure 2. Overview of ShapeMoiré. The top part depicts the network’s overall architecture. The green and blue parts denote the baseline and ShapeMoiré components, respectively. and The red line flow represents ShapeMoiré during inference. The bottom part shows network layers: (a) Convolution-Layer (consistent in both training and testing), and (b) and (c) illustrate ShapeConv during training and inference, respectively. For both ShapeConv and Shape-Architecture, our method introduces no parameters and negligible computational cost during inference.